A Gallup poll found 57% of people were pro-choice on abortion.
1. Find the probability that of the next 40 randomly sampled people that 25 (x = 25) are pro-choice
2. Find the probability that of the next 30 randomly sampled people that less than 15 (x ≤ 15) are pro-choice
3. Find the probability that of the next 50 randomly sampled people that between 20 and 30 (20 ≤ x ≤ 30)are pro-choice.
Answer #1

p=054 2=*-* 1: n=40 P1x225) - H0 (0.57² (0.43) = 10.1008] 2. n=30 M = m p=30x0.57 = 17.1 2=X-M. a=5hp(16) Boxo. 59 80.43 22.7